# BALROG: Benchmarking Agentic LLM and VLM Reasoning On Games

### v1

@@ Abstract
Large Language Models (LLMs) and Vision Language Models (VLMs) possess extensive knowledge and exhibit promising reasoning abilities, however, they still struggle to perform well in complex, dynamic environments. Real-world tasks require handling intricate interactions, advanced spatial reasoning, long-term planning, and continuous exploration of new strategies—areas in which we lack effective methodologies for comprehensively evaluating these capabilities. To address this gap, we introduce BALROG, a novel benchmark designed to assess the agentic capabilities of LLMs and VLMs through a diverse set of challenging games. Our benchmark incorporates a range of existing reinforcement learning environments with varying levels of difficulty, including tasks that are solvable by non-expert humans in seconds to extremely challenging ones that may take years to master (e.g., the NetHack Learning Environment). We devise fine-grained metrics to measure performance and conduct an extensive evaluation of several popular open-source and closed-source LLMs and VLMs. Our findings indicate that while current models achieve partial success in the easier games, they struggle significantly with more challenging tasks. Notably, we observe severe deficiencies in vision-based decision-making, as models perform worse when visual representations of the environments are provided. We release BALROG as an open and user-friendly benchmark to facilitate future research and development in the agentic community.
@@ 1 Introduction
Current agentic benchmarks evaluate LLM performance in settings that involve no more than a few dozen rounds of interaction between a model and an environment, e.g., solving simple office tasks (Wang et al., 2024), navigating the Internet  (Zhou et al., 2023), and resolving GitHub issues (Jimenez et al., 2023). New agentic prompting frameworks and improvements to short-horizon reasoning via LLMs like OpenAI o1 have led to dramatic and fast-paced gains in state-of-the-art performance on these benchmarks (OpenAI, 2024b; Wang et al., 2023; Fernando et al., 2023; Hu et al., 2024). However, many realistic tasks require orders of magnitude more interactions (Pignatiello et al., 2020; Wansink and Sobal, 2007).
@@
In this paper, we argue that the next frontier for language and vision-language model capabilities lies in long-horizon reasoning and decision-making. To that end, we propose BALROG: Benchmarking Agentic LLM/VLM Reasoning On Games. BALROG is a benchmark and framework that aggregates a diverse set of complex reinforcement learning game environments into a unified testbed for research on long-context LLMs. Games have historically served as highly effective metrics for evaluating progress in deep reinforcement learning research (Bellemare et al., 2013; Silver et al., 2018; Schrittwieser et al., 2020; Vinyals et al., 2019). By aggregating many different game environments into a single evaluation, we look to spur progress on developing truly generalist agents that can meaningfully address embodied, real world tasks. Specifically BALROG enables seamless running of LLM and VLM agents on BabyAI, Crafter, TextWorld, Baba Is AI, MiniHack, and NetHack (Chevalier-Boisvert et al., 2019; Hafner, 2021; Côté et al., 2019; Cloos et al., 2024; Samvelyan et al., 2021; Küttler et al., 2020). These environments have lightweight simulators, ensuring that the benchmark is affordable for the research community. Furthermore, while all of these games are long-horizon, they span a broad range of difficulty levels, from tasks where we see fair zero-shot performance by state-of-the-art long-context models (BabyAI) to those where even specialized neural models trained on billions of in-domain datapoints make very limited progress (NetHack) (Piterbarg et al., 2024; Klissarov et al., 2023; Wołczyk et al., 2024). BALROG is difficult to solve through simple memorization – all of the environments used in the benchmark are procedurally generated, and encountering the same instance of an environment twice is unlikely.
@@ 3.1 Evaluation Setting
To address cases where the LLMs/VLMs output hallucinated or invalid actions, BALROG provides feedback to the agent indicating the action’s invalidity, it then executes a default fallback action (such as a “do-nothing” action or a standard move like “north”), and logs the occurrence for trajectory statistics. This ensures that the interaction remains continuous and robust while enabling users to analyze the context and frequency of such errors in post-evaluation analysis.
@@ 3.2 Observations
In the initial prompt, the agent is introduced to the game rules and provided with a list of available actions, each accompanied by a brief description. To prevent model overspecialization, we design a general prompt that is not fine-tuned to any specific LLM. Subsequent prompts present the observation-action history in a chat-based format. The game rules and observations are conveyed from the perspective of the “user”, while prior actions are attributed to the “assistant” or “model” role, depending on the type of model used. This structure mirrors the standard format used for fine-tuning instruction-following LLMs. Detailed examples of game observations are included in the appendices.
@@
Language Only Format Observations are expressed as natural language descriptions of the environment’s state (e.g., “a wall 5 steps ahead, a wall 2 steps to the left…”). For environments without native textual representations, we either generate descriptions using open-source language wrappers (BabyAI (Carta et al., 2023), Crafter (Wu et al., 2023), NetHack, and MiniHack (Goodger et al., 2023)) or develop a custom wrapper ourselves (Baba is AI, see Appendix 5)
@@ 3.3 Models
We evaluate a range of popular closed-source and open-source models, including Gemini-1.5-Flash and Gemini-1.5-Pro (Reid et al., 2024), GPT-4o-mini (2024-07-18 release) and GPT-4o (2024-05-13 release) (Achiam et al., 2023; OpenAI, 2024a), Claude 3.5 Sonnet (Anthropic, 2024), as well as Llama 3.1 instruct (8B and 70B) (Dubey et al., 2024) and Llama 3.2 instruct (1B, 3B, 11B and 90B) (MetaAI, 2024). Additionally, we test o1-mini (2024-09-12 release) and o1-preview (2024-09-12 release) (OpenAI, 2024b) exclusively on the NetHack environment due to budget constraints.
@@ 4 Results
In Figure 2, we present the results of our experiments using the BALROG evaluation script for both language-only and vision-language formats. Most leading models demonstrate fair average progression on BabyAI, Crafter, and Baba Is AI, with GPT-4o performing best. Interestingly, the open-source Llama 3.1 70B and Llama 3.2 90B models achieve the highest results on the Baba Is AI language-only format, narrowly surpassing GPT-4o and Claude 3.5 Sonnet. In TextWorld, GPT-4o and Claude 3.5 Sonnet lead, while Gemini models fail to complete any tasks, being flagged as ‘unsafe’ by the Google Gemini API, despite the prompts containing no actual safety concerns. The MiniHack suite proves very challenging for all models, especially the quest and boxoban tasks, which were never solved by any model. Finally, all models flat line with NetHack, with the best-performing model, o1-preview, achieving a meager 1.5% average game progression.
Table LABEL:llm_table summarizes the aggregated results across all environments in the language-only format. Overall, GPT-4o is the best-performing model, with an average progression of 31.62%, followed closely by Claude 3.5 Sonnet and Llama 3.1 70B. Gemini-1.5-Pro lags behind the other large models, partly due to its 0% performance on TextWorld. However, results differ for the vision-language format, as shown in Table LABEL:vlm_table. Here, we observe that both GPT-4o and Llama 3.2 exhibit a decline in performance when image observations are included, likely due to confusion arising from the added visual input. In contrast, Gemini-1.5-Pro and Claude 3.5 Sonnet especially, maintain consistent performance across both formats. This suggests that current multimodal Transformer architectures are still better equipped at handling textual information than visual input, a topic we explore further in Section 6. Additionally, Llama 3.1 70B outperforms the larger and more recent Llama 3.2 90B in the language-only format, suggesting that the introduction of visual processing in the latter may have negatively impacted its linguistic and reasoning capabilities. We show more detailed results for each environment in their appendices.
@@
llama-3.1-8b-it 14.14 $\pm$ 1.51
llama-3.2-11B-it 13.54 $\pm$ 1.05
gemini-1.5-flash 9.73 $\pm$ 0.77
llama-3.2-3B-it 8.47 $\pm$ 1.12
llama-3.2-1B-it 6.32 $\pm$ 1.00
@@ 5 Related Work
In this work, we study agentic skills in the context of video games, as they offer challenges well-tailored for human players and test skills that are useful for embodied agents. Previously, some related works employed games to benchmark LLMs (Liu et al., 2023b; Todd et al., 2024; Wu et al., 2023), highlighting their emphasis on problem-solving, spatial reasoning, and well-defined rules and objectives. Some of these benchmarks, however, are already reaching saturation, with environments like Crafter being the most challenging in their suite. In contrast, BALROG fills an important gap by providing a wide range of games at varying difficulties—including the NetHack Learning Environment (Küttler et al., 2020), which takes humans years to master, and where zero-shot LLMs struggle greatly, as also seen in prior work (Jeurissen et al., 2024). These tasks represent a rich and granular testbed for evaluating agentic foundation models, pushing decision-making evaluations of LLMs/VLMs to the very limit of their context lengths. Other environments such as MineDojo (Fan et al., 2022) and MineRL (Guss et al., 2019) also present open-ended challenges for agentic capabilities, their steep computational requirements and reliance on multimodal inputs make them less practical for accessible, large-scale benchmarks.
@@ 6 Open Research Problems
Beyond simply prompting LLMs and VLMs to directly predict the next action of game-play, BALROG also supports the study of more advanced reasoning techniques like chain-of-thought (Wei et al., 2022b), self-refinement (Madaan et al., 2024), and basic planning. These methods have been demonstrated to improve model performance on shorter-context problems. We believe them to be an exciting direction for future work on long-context reasoning and decision-making. For example, model performance on the tasks in BALROG might be improved by integrating multi-agent collaboration (Chang, 2023; Khan et al., 2024; Yao et al., 2024) and tool usage (Shen et al., 2024; Ruan et al., 2023; Schick et al., 2024; Qin et al., 2023) in decision-making. Additionally, incorporating memory mechanisms or reinforcement learning techniques could help bridge the “knowing-doing” gap, enabling models to apply their knowledge effectively in practical, long-horizon tasks. Finally, experimenting with open-ended self-improvement loops (Wang et al., 2023; Hu et al., 2024) could lead to more adaptive and general agents (Team et al., 2023; Hughes et al., 2024), offering a pathway toward truly autonomous systems.
@@
Despite their potential, our benchmark shows significant variability in VLM performance. While some models, like Llama 3.2, struggle to integrate visual information into coherent decision-making, others—most notably Sonnet 3.5—demonstrate stronger performance in VLM mode. This disparity highlights significant variability in VLM capabilities, which may stem from differences in training objectives and datasets. For example, Sonnet 3.5’s superior performance can be attributed in part to its training on tasks involving computer usage (Anthropic, 2024), which inherently require integrating visual and textual inputs for action-based reasoning.
@@ Appendix A Baby AI
BabyAI [Chevalier-Boisvert et al., 2019] is a research platform designed to study grounded language learning and instruction following in artificial agents. It consists of a suite of 2D grid world environments with increasing levels of complexity. In these environments, an agent navigates through rooms and interacts with various objects like doors, keys, balls, and boxes of different colors. The agent receives natural language instructions, called “missions”, which describe tasks it needs to complete, such as picking up specific objects or navigating to certain locations. Many existing works on decision-making have studied model performance on this environment [Reed et al., 2022, Li et al., 2022]. We use it as a historically relevant environment that we expect to be relatively easy to solve.
@@ A.1 BabyAI-Text
We evaluate the agents on 5 tasks introduced in BabyAI-Text [Carta et al., 2023], which provides a description of each observation instead of a symbolic representation. A textual description consists of a list of template descriptions with the following structure:
@@ A.2 BabyAI Results
llama-3.2-11B-it 32.80 $\pm$ 4.20
llama-3.1-8B-it 30.00 $\pm$ 6.48
gemini-1.5-flash 25.60 $\pm$ 3.90
llama-3.2-3B-it 10.00 $\pm$ 4.24
llama-3.2-1B-it 6.00 $\pm$ 3.36
@@ Appendix B Crafter
Crafter [Hafner, 2021] is an open-source 2D survival game designed specifically for research on strong generalization, deep exploration, and long-term reasoning in reinforcement learning. It is a Minecraft-inspired, procedurally generated environment that combines resource gathering, crafting, and combat elements. Additionally, the game includes a comprehensive set of tasks and achievements, enabling researchers to evaluate agent performance across multiple objectives and time scales. To enable interaction with language models we use the same language wrapper as proposed in  Wu et al. [2023].
@@ B.1 Crafter Results
We provide Crafter results for LLM and VLM format in Tables 7 and 7, standard errors are computed using 10 seeds. GPT4o leads in language-only mode, and Gemini-1.5-Pro leads in vision-language mode. Surprisingly, Llama 3.2 90B performance decreases very sharply when images are added, getting worse average progress than its smaller 11B model.

### v2

@@ Abstract
Large Language Models (LLMs) and Vision Language Models (VLMs) possess extensive knowledge and exhibit promising reasoning abilities, however, they still struggle to perform well in complex, dynamic environments. Real-world tasks require handling intricate interactions, advanced spatial reasoning, long-term planning, and continuous exploration of new strategies-areas in which we lack effective methodologies for comprehensively evaluating these capabilities. To address this gap, we introduce BALROG, a novel benchmark designed to assess the agentic capabilities of LLMs and VLMs through a diverse set of challenging games. Our benchmark incorporates a range of existing reinforcement learning environments with varying levels of difficulty, including tasks that are solvable by non-expert humans in seconds to extremely challenging ones that may take years to master (e.g., the NetHack Learning Environment). We devise fine-grained metrics to measure performance and conduct an extensive evaluation of several popular open-source and closed-source LLMs and VLMs. Our findings indicate that while current models achieve partial success in the easier games, they struggle significantly with more challenging tasks. Notably, we observe severe deficiencies in vision-based decision-making, as several models perform worse when visual representations of the environments are provided. We release BALROG as an open and user-friendly benchmark to facilitate future research and development in the agentic community. Code and Leaderboard at balrogai.com.
@@ 1 Introduction
Current agentic benchmarks evaluate LLM performance in settings that involve no more than a few dozen rounds of interaction between a model and an environment, e.g., solving simple office tasks (Wang et al., 2024), navigating the Internet  (Zhou et al., 2023), and resolving GitHub issues (Jimenez et al., 2023). New agentic prompting frameworks and improvements to short-horizon reasoning via LLMs like OpenAI o1 have led to dramatic and fast-paced gains in state-of-the-art performance on these benchmarks (OpenAI, 2024b; Wang et al., 2023; Fernando et al., 2023; Hu et al., 2024). However, many realistic tasks require orders of magnitude more interactions (Pignatiello et al., 2020; Wansink & Sobal, 2007).
@@
In this paper, we argue that the next frontier for language and vision-language model capabilities lies in long-horizon reasoning and decision-making. To that end, we propose BALROG: Benchmarking Agentic LLM and VLM Reasoning On Games. BALROG is a benchmark and framework that aggregates a diverse set of complex reinforcement learning game environments into a unified testbed for research on long-context LLMs. Games have historically served as highly effective metrics for evaluating progress in deep reinforcement learning research (Bellemare et al., 2013; Silver et al., 2018; Schrittwieser et al., 2020; Vinyals et al., 2019). By aggregating many different game environments into a single evaluation, we look to spur progress on developing truly generalist agents that can meaningfully address embodied, real world tasks. Specifically BALROG enables seamless running of LLM and VLM agents on BabyAI, Crafter, TextWorld, Baba Is AI, MiniHack, and NetHack (Chevalier-Boisvert et al., 2019; Hafner, 2021; Côté et al., 2019; Cloos et al., 2024; Samvelyan et al., 2021; Küttler et al., 2020). These environments have lightweight simulators, ensuring that the benchmark is affordable for the research community. Furthermore, while all of these games are long-horizon, they span a broad range of difficulty levels, from tasks where we see fair zero-shot performance by state-of-the-art long-context models (BabyAI) to those where even specialized neural models trained on billions of in-domain datapoints make very limited progress (NetHack) (Piterbarg et al., 2024; Klissarov et al., 2023; Wołczyk et al., 2024). BALROG is difficult to solve through simple memorization – all of the environments used in the benchmark are procedurally generated, and encountering the same instance of an environment twice is unlikely.
@@ 3.1 Evaluation Setting
To address cases where the LLMs/VLMs output hallucinated or invalid actions, BALROG provides feedback to the agent indicating the action’s invalidity, it then executes a default fallback action (such as a ”do-nothing” action or a standard move like ”north”), and logs the occurrence for trajectory statistics. This ensures that the interaction remains continuous and robust while enabling users to analyze the context and frequency of such errors in post-evaluation analysis.
@@ 3.2 Observations
In the initial prompt, the agent is introduced to the game rules and provided with a list of available actions, each accompanied by a brief description. To prevent model overspecialization, we design a general prompt that is not fine-tuned to any specific LLM. Subsequent prompts present the observation-action history in a chat-based format. The game rules and observations are conveyed from the perspective of the ”user”, while prior actions are attributed to the ”assistant” or ”model” role, depending on the type of model used. This structure mirrors the standard format used for fine-tuning instruction-following LLMs. Detailed examples of game observations are included in the appendices.
@@
Language Only Format Observations are expressed as natural language descriptions of the environment’s state (e.g., ”a wall 5 steps ahead, a wall 2 steps to the left…”). For environments without native textual representations, we either generate descriptions using open-source language wrappers (BabyAI (Carta et al., 2023), Crafter (Wu et al., 2023), NetHack, and MiniHack (Goodger et al., 2023)) or develop a custom wrapper ourselves (Baba is AI, see Appendix 6)
@@ 3.3 Models
We evaluate a range of popular closed-source and open-source models, including Gemini-1.5-Flash and Gemini-1.5-Pro (Reid et al., 2024), GPT-4o-mini (2024-07-18 release) and GPT-4o (2024-05-13 release) (Achiam et al., 2023; OpenAI, 2024a), Claude 3.5 Sonnet and Claude 3.5 Haiku (2024-10-22 releases) (Anthropic, 2024), as well as Llama 3.1 instruct (8B and 70B) (Dubey et al., 2024) and Llama 3.2 instruct (1B, 3B, 11B and 90B) (MetaAI, 2024). Additionally, we test o1-preview (2024-09-12 release) (OpenAI, 2024b) exclusively on the NetHack environment due to budget constraints.
@@ 4 Results
In Figure 2, we present the results of our experiments using the BALROG evaluation script for both language-only and vision-language formats. Most leading models demonstrate fair average progression on BabyAI, Crafter, and Baba Is AI, with GPT-4o and Claude 3.5 Sonnet performing best. Interestingly, the open-source Llama 3.1 70B and Llama 3.2 90B models achieve the highest results on the Baba Is AI language-only format, narrowly surpassing GPT-4o and Claude 3.5 Sonnet. In TextWorld, GPT-4o and Claude 3.5 Sonnet lead, while Gemini models fail to complete any tasks, being flagged as ’unsafe’ by the Google Gemini API, despite the prompts containing no actual safety concerns. The MiniHack suite proves very challenging for all models, especially the quest and boxoban tasks, which were never solved by any model. Finally, all models flat line with NetHack, with the best-performing model, o1-preview, achieving a meager 1.5% average game progression.
Table LABEL:llm_table summarizes the aggregated results across all environments in the language-only format. Overall, Claude 3.5 Sonnet is the best-performing model, with an average progression of 32.64%, followed closely by GPT-4o and Llama 3.1 70B a few points behind. Gemini-1.5-Pro lags behind the other large models, partly due to its 0% performance on TextWorld. However, results differ for the vision-language format, as shown in Table LABEL:vlm_table. Here, we observe that both GPT-4o and Llama 3.2 exhibit a decline in performance when image observations are included, likely due to confusion arising from the added visual input. In contrast, Gemini-1.5-Pro and Claude 3.5 Sonnet especially, maintain consistent performance across both formats. This suggests that current multimodal Transformer architectures are still better equipped at handling textual information than visual input, a topic we explore further in Section 6. We show more detailed results for each environment in their appendices.
@@
llama-3.2-11B-it 16.82 $\pm$ 1.47
llama-3.1-8b-it 15.14 $\pm$ 1.55
gemini-1.5-flash 14.63 $\pm$ 1.37
llama-3.2-3B-it 10.13 $\pm$ 1.28
llama-3.2-1B-it 6.65 $\pm$ 1.04
@@ 5 Related Work
In this work, we study agentic skills in the context of video games, as they offer challenges well-tailored for human players and test skills that are useful for embodied agents. Previously, some related works employed games to benchmark LLMs (Liu et al., 2023b; Todd et al., 2024; Wu et al., 2023; Ruoss et al., 2024), highlighting their emphasis on problem-solving, spatial reasoning, and well-defined rules and objectives. Some of these benchmarks, however, are already reaching saturation, with environments like Crafter being the most challenging in their suite. In contrast, BALROG fills an important gap by providing a wide range of games at varying difficulties-including the NetHack Learning Environment (Küttler et al., 2020), which takes humans years to master, and where zero-shot LLMs struggle greatly, as also seen in prior work (Jeurissen et al., 2024). These tasks represent a rich and granular testbed for evaluating agentic foundation models, pushing decision-making evaluations of LLMs/VLMs to the very limit of their context lengths. Other environments such as MineDojo (Fan et al., 2022) and MineRL (Guss et al., 2019) also present open-ended challenges for agentic capabilities, their steep computational requirements and reliance on multimodal inputs make them less practical for accessible, large-scale benchmarks.
@@ 6 Open Research Problems
Beyond simply prompting LLMs and VLMs to directly predict the next action of game-play, BALROG also supports the study of more advanced reasoning techniques like chain-of-thought (Wei et al., 2022b), self-refinement (Madaan et al., 2024), and basic planning. These methods have been demonstrated to improve model performance on shorter-context problems. We believe them to be an exciting direction for future work on long-context reasoning and decision-making. For example, model performance on the tasks in BALROG might be improved by integrating multi-agent collaboration (Chang, 2023; Khan et al., 2024; Yao et al., 2024) and tool usage (Shen et al., 2024; Ruan et al., 2023; Schick et al., 2024; Qin et al., 2023) in decision-making. Additionally, incorporating memory mechanisms or reinforcement learning techniques could help bridge the ”knowing-doing” gap, enabling models to apply their knowledge effectively in practical, long-horizon tasks. Finally, experimenting with open-ended self-improvement loops (Wang et al., 2023; Hu et al., 2024) could lead to more adaptive and general agents (Team et al., 2023; Hughes et al., 2024), offering a pathway toward truly autonomous systems.
@@
Despite their potential, our benchmark shows significant variability in VLM performance. While some models, like Llama 3.2, struggle to integrate visual information into coherent decision-making, others-most notably Sonnet 3.5-demonstrate stronger performance in VLM mode. This disparity highlights significant variability in VLM capabilities, which may stem from differences in training objectives and datasets. For example, Sonnet 3.5’s superior performance can be attributed in part to its training on tasks involving computer usage (Anthropic, 2024), which inherently require integrating visual and textual inputs for action-based reasoning.
@@ 7 Conclusion
Acknowledgments
We thank the Gemini Academic Program and Divy Thakkar for their support. We further thank Roberta Raileanu, Pierluca d’Oro, Mikayel Samvelyan, Sam Devlin, Eric Hambro, and Heinrich Kttler for the insightful discussions.
@@ Appendix A Baby AI
BabyAI (Chevalier-Boisvert et al., 2019) is a research platform designed to study grounded language learning and instruction following in artificial agents. It consists of a suite of 2D grid world environments with increasing levels of complexity. In these environments, an agent navigates through rooms and interacts with various objects like doors, keys, balls, and boxes of different colors. The agent receives natural language instructions, called “missions”, which describe tasks it needs to complete, such as picking up specific objects or navigating to certain locations. Many existing works on decision-making have studied model performance on this environment (Reed et al., 2022; Li et al., 2022). We use it as a historically relevant environment that we expect to be relatively easy to solve.
@@ A.1 BabyAI-Text
We evaluate the agents on 5 tasks introduced in BabyAI-Text (Carta et al., 2023), which provides a description of each observation instead of a symbolic representation. A textual description consists of a list of template descriptions with the following structure:
@@ A.2 BabyAI Results
gemini-1.5-flash 50.00 $\pm$ 7.07
llama-3.2-11b-it 50.00 $\pm$ 7.07
llama-3.1-8b-it 36.00 $\pm$ 6.79
llama-3.2-3b-it 20.00 $\pm$ 5.66
llama-3.2-1b-it 8.00 $\pm$ 3.84
@@ Appendix B Crafter
Crafter (Hafner, 2021) is an open-source 2D survival game designed specifically for research on strong generalization, deep exploration, and long-term reasoning in reinforcement learning. It is a Minecraft-inspired, procedurally generated environment that combines resource gathering, crafting, and combat elements. Additionally, the game includes a comprehensive set of tasks and achievements, enabling researchers to evaluate agent performance across multiple objectives and time scales. To enable interaction with language models we use the same language wrapper as proposed in  Wu et al. (2023).
@@ B.1 Crafter Results
We provide Crafter results for LLM and VLM format in Tables LABEL:crafter_LLM and LABEL:crafter_VLM, standard errors are computed using 10 seeds. GPT4o leads in language-only mode, and Gemini-1.5-Pro leads in vision-language mode. Surprisingly, Llama 3.2 90B performance decreases very sharply when images are added, getting worse average progress than its smaller 11B model.
@@ 22. Collect Diamond
B.3 Analysis of VLM Performance in the Crafter Environment
To investigate the agentic capabilities of VLMs in visually complex environments, we focused on the Crafter environment, a simplified 2D version of Minecraft. The environment uses tiled 2D textures to render scenes, which may be out-of-distribution for many VLMs. To address this potential limitation, we introduced an augmented version of the environment where scenes are rendered in 3D using Minecraft’s 3D models and textures. This approach leverages data types more likely to be present in the training datasets of VLMs, given the abundance of interleaved text and images about Minecraft on the web.
Results
Surprisingly, the switch to 3D rendering did not improve VLM performance. In some cases, performance on the 3D-rendered environment was slightly worse than on the original 2D textures. This result challenges the hypothesis that familiarity with Minecraft-like 3D data in training datasets would lead to better performance. Instead, it suggests that VLMs may struggle with agentic tasks for reasons unrelated to the type of visual texture or rendering style.
Discussion
Our findings indicate that current VLM models may not be well-suited for complex vision-based reasoning tasks, even when provided with familiar visual contexts. The lack of improvement with 3D rendering suggests that factors other than texture familiarity, such as limitations in spatial reasoning or task-specific adaptation, may play a more significant role in their underperformance. While these preliminary results are insightful, further experimentation is necessary to draw stronger conclusions.
Figure 4: Crafter’s example of 3d scene visualization with Minecraft 3d models and textures.
Table 8: Average Progress of VLMs in Crafter Across Rendering Types
Model 3d textures(%) 2d textures(%)
claude-3.5-sonnet 25.45 $\pm$ 4.18 37.27 $\pm$ 3.14
gemini-1.5-pro 30.45 $\pm$ 3.08 33.50 $\pm$ 2.07
gpt-4o 24.55 $\pm$ 3.65 26.81 $\pm$ 3.74
gemini-1.5-flash 19.09 $\pm$ 2.78 20.70 $\pm$ 4.43
gpt-4o-mini 17.27 $\pm$ 2.38 19.91 $\pm$ 3.13
llama-3.2-11B-it 15.90 $\pm$ 1.15 23.63 $\pm$ 1.48
llama-3.2-90B-it 12.27 $\pm$ 1.44 10.00 $\pm$ 1.13